OpenStudy (anonymous):

What is the standard form of the equation of a circle with its center at (2, -3) and passing through the point (-2, 0)? (x - 2)2^ + (y + 3)2^ = 5 (x + 2)2^ + (y - 3)2^ = 25 (x - 2)2^ + (y + 3)2^ = 25 (x - 2)2^ - (y + 3)2^ = 5

OpenStudy (anonymous):

@Sasarapony123

OpenStudy (anonymous):

@shyboo1998

OpenStudy (anonymous):

@K_V8

OpenStudy (anonymous):

ok so whats the question

OpenStudy (anonymous):

^^^

OpenStudy (anonymous):

@Sasarapony123 HELP

OpenStudy (anonymous):

uh my math isn't that advanced yet sorry

OpenStudy (anonymous):

ill try my best though

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

I only have 15 mins left of class

OpenStudy (anonymous):

have to go

OpenStudy (anonymous):

u have to go?

OpenStudy (anonymous):

A circle is described by the equation (x−12)2+(y−32)2=49. What are the cordinates for the center of the circle and the length of the radius?

OpenStudy (anonymous):

(-12,-32), 23units (-12,-32), 49units (12,32), 49units (12,32), 23units

OpenStudy (anonymous):

Try it and ill be back tomorrow

OpenStudy (anonymous):

the ????????? are subtraction

OpenStudy (anonymous):

What is the general form of the equation for the given circle centered at O(0, 0)? x2 + y2 + 41 = 0 x2 + y2 − 41 = 0 x2 + y2 + x + y − 41 = 0 x2 + y2 + x − y − 41 = 0

OpenStudy (anonymous):

What is the general form of the equation of a circle with center at (a, b) and radius of length m? x2 + y2 − 2ax − 2by + (a2 + b2 − m2) = 0 x2 + y2 + 2ax + 2by + (a2 + b2 − m2) = 0 x2 + y2 − 2ax − 2by + (a + b − m2) = 0 x2 + y2 + 2ax + 2by + a2 + b2 = -m2